40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13137     Entry time: Mon Jul 24 12:00:21 2017
Author: gautam 
Type: Update 
Category: PSL 
Subject: PSL NPRO mysteriously shut off 

Summary:

At around 10:30AM today morning, the PSL mysteriously shut off. Steve and I confirmed that the NPRO controller had the RED "OFF" LED lit up. It is unknown why this happened. We manually turned the NPRO back on and hte PMC has been stably locked for the last hour or so.

Details:

There are so many changes to lab hardware/software that have been happening recently, it's not entirely clear to me what exactly was the problem here. But here are the observations:

  1. Yesterday, when I came into the lab, the MC REFL trace on the wall StripTool was 0 for the full 8 hour history - since we don't have data records, I can't go back further than this. I remember the PMC TRANS and REFL cameras looked normal, but there was no MC REFL spot on the CCD monitors. This is consistent with the PSL operating normally, the PMC being locked, and the PSL shutter being closed. Isn't the emergency vacuum interlock also responsible for automatically closing the PSL shutter? Perhaps if the turbo controller failure happened prior to Jamie/me coming in yesterday, maybe this was just the interlock doing its job. On Friday evening, the PSL shutter was certainly open and the MC REFL spot was visible on the camera. I also confirmed with Jamie that he didn't close the shutter.
  2. Attachment #1 shows the wall StripTool traces from earlier this morning. It looks like ~7.40AM, the MC REFL level went back up. Steve says he didn't manually open the shutter, and in any case, this was before the turbo pump controller failure was diagnosed. So why did the shutter open again
  3. When I came in at ~10AM, the CCD monitor showed that the PMC was locked, and the MC REFL spot was visible. 
  4. Also on attachment #1, there is a ~10min dip in the MC REFL level. This corresponds to ~10:30AM this morning. Both Steve and I were sitting in the control room at this time. We noticed that the PMC TRANS and REFL CCDs were dark. When we went in to check on the laser, we saw that it was indeed off. There was no one inside the lab area at this time to our knowledge, and as far as I know, the only direct emergency shutoff for the PSL is on the North-West corner of the PSL enclosure. So it is unclear why the laser just suddenly went off.

Steve says that this kind of behaviour is characteristic of a power glitch/surge, but nothing else seems to have been affected (I confirmed that the X and Y end lasers are ON). 

Attachment 1: IMG_7454.JPG  1.622 MB  | Hide | Hide all
IMG_7454.JPG
ELOG V3.1.3-